OpenStudy (anonymous):

What is the exact value of: The square root of 27/64?

OpenStudy (anonymous):

Take the square root of the top and the bottom. I cant say what is the exact value as square root of 27 is irrational.

OpenStudy (anonymous):

ok actually you are looking for \[\sqrt{\frac{3^3}{8^2}}\]

OpenStudy (anonymous):

\[\sqrt{\frac{3^3}{8^2}}=\frac{\sqrt{3^3}}{\sqrt{8^2}}\] \[=\frac{3\sqrt{3}}{8}\]

OpenStudy (anonymous):

(3*sqrt(3))/8

OpenStudy (anonymous):

Thanks guys :)